# Using Language and Road Manuals to Inform Map Reconstruction for Autonomous Driving

### v1

@@ I Introduction
A key limitation of these works is that they generally only use road geometry and the road-type labels (e.g:, highway=residential) from OSM while ignoring the other semantic and geometric information available for each road. Lane-level cues, such as the number of lanes and lane width, are also overlooked, even though they could aid lane-topology prediction in scenarios with camera occlusion, unmarked roads, or complex road intersections.
@@ II-B Online Lane-Topology Prediction
SD Maps like OSM provide the advantage of being available open-source and for a major portion of the world. OSM also provides road metadata such as the number of lanes, service classification of the road, road direction, and speed limit, etc., which has crucial semantic information captured in natural language. This information can help disambiguate road functionality and infer lane connectivity. One key detail that is very inconsistently tagged in OSM is the lane width of a road. Lane widths can help deduce the spatial layout of multi-lane roads even in the absence of clear lane markings. Diwanji et al. [4] show that the standardized road design manuals, such as the Highway Design Manual [14], provide the lane widths for different road types which can be used for lane-topology prediction.
@@ II-C Case Study: Real-world applicability of Online Mapping Methods
These works report that while the StreamMapNet-based model performs very closely to the HD Map-based model, the model that uses lane-centerline-predicting variant of MapTRv2 even outperforms the HD Map-based model. Such results demonstrate the viability of online lane-topology prediction methods as alternatives to HD Maps for downstream Autonomous Vehicles tasks.
@@ IV-B Evaluation Metrics
We also include a fifth metric, OpenLane-V2 Score (OLS) which is the weighted average of all four metrics.
@@ IV-C1 Model Training Configurations
  • F0: We freeze the entire model, except the 2-layer MLP for transforming the text embeddings, for the second round of training

### v2

@@ I Introduction
A key limitation of these works is that they generally only use road geometry and the road-type labels (e.g, highway=residential) from OSM while ignoring the other semantic and geometric information available for each road. Lane-level cues, such as the number of lanes and lane width, are also overlooked, even though they could aid lane-topology prediction in scenarios with camera occlusion, unmarked roads, or complex road intersections.
@@ II-B Online Lane-Topology Prediction
SD Maps like OSM provide the advantage of being available open-source and for a major portion of the world. OSM also provides road metadata such as the number of lanes, service classification of the road, road direction, and speed limit, etc., which has crucial semantic information captured in natural language. This information can help disambiguate road functionality and infer lane connectivity. One key detail that is very inconsistently tagged in OSM is the lane width of a road. Lane widths can help deduce the spatial layout of multi-lane roads even in the absence of clear lane markings. Diwanji et al. [4] show that the standardized road design manuals, such as the Highway Design Manual [14], provide the lane widths for different road types, which can be used for lane-topology prediction.
@@ II-C Case Study: Real-world applicability of Online Mapping Methods
These works report that while the StreamMapNet-based model performs very closely to the HD Map-based model, the model that uses the lane-centerline-predicting variant of MapTRv2 even outperforms the HD Map-based model. Such results demonstrate the viability of online lane-topology prediction methods as alternatives to HD Maps for downstream Autonomous Vehicles tasks.
@@ IV-B Evaluation Metrics
We also include a fifth metric, OpenLane-V2 Score (OLS), which is the weighted average of all four metrics.
@@ IV-C1 Model Training Configurations
  • F0: We freeze the entire model, except the 2-layer MLP for transforming the text embeddings, for the second round of training.